40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 454     Entry time: Tue Jan 30 09:19:09 2018
Author: Gabriele 
Type: Optics 
Category: Configuration 
Subject: Setup for 50mm disks 

I realigned all optical levers to measure the 50mm disks. In brief, I moved the input 2" mirrors, the in-vacuum 2" mirrors and the PZT mirrors so that the beam hits the 50mm sample and gets back into the QPD. Re-aligned everything to the horizontal reference using water.

ELOG V3.1.3-